OpenStudy (anonymous):

One pipe can fill the tank in 5 hours. A second can fill it in 3 hours. How long would it take both pipes to fill the tank 3/4 of the way full?

hero (hero):

Hint: \[\frac{x}{3} + \frac{x}{5} = \frac{3}{4}\] Solve for x

OpenStudy (anonymous):

ok thanks